OpenStudy (anonymous):

solve: ( 1/(e^x) ) = 5

OpenStudy (anonymous):

This is tricky one you will have to use natural log.

OpenStudy (anonymous):

\[\frac{1}{e^x}=5\]\[1=5*e^x\]\[\frac{1}{5}=e^x\]

OpenStudy (zzr0ck3r):

1/e^(x) = 5 e^(x) = 1/5 ln(e^(x)) = ln(1/5) x = ln(1/5) = ln(1) - ln(5) = 0 - ln(5) = -ln(5)

OpenStudy (anonymous):

from there you simply natural log both sides.

OpenStudy (anonymous):

\[e^{-x}=5\] does this look better? or maybe \[e^{x}= \frac{1}{5}\]

OpenStudy (anonymous):

The natural log will cancel the e and bring down the x

OpenStudy (anonymous):

You understand?

OpenStudy (anonymous):

so the answer would be ln(1/5) ?

OpenStudy (lgbasallote):

haha so many answers

OpenStudy (anonymous):

Yes that is the answer. You can simplify it however you want. I would just leave it as that.

OpenStudy (anonymous):

okay thank you. that makes sense!

OpenStudy (zzr0ck3r):

notice that 1/5 = 5^(-1) and a property of logs is a*log(b) = log(b^a) so log(1/5) = log(5^(-1)) = -1*log(5) this is the great thing about logs, you can really "play around" with them. notice how i got to the same thing two compelety different ways.